OpenStudy (anonymous):

A motorboat heads due east at 10.0 m/s. The river has a current of 6 m/s due north. What is the magnitude of the resultant velocity of the boat?

OpenStudy (anonymous):

 \[R^2 = 10^2 +6^2 . Resultant velocity =11.6 m/s\]